what is focus and directix of parabola y^2=4(y-1)
Answer by lwsshak3(11628)

what is focus and directix of parabola y^2=4(y-1)
I think you meant to write the equation as y^2=4(x-1)
This is an equation of a parabola that opens right.
Its standard form: (y-k)^2=4p(x-h), (h,k)=(x,y) coordinates of the vertex
For given parabola:
vertex: (1,0)
axis of symmetry: y=0 or x-axis
4p=4
p=1
focus: (2,0) (p-distance right of vertex on the axis of symmetry)
directrix: x=0 or y-axis (p-distance left of vertex on the axis of symmetry)